# Self Iterative Label Refinement via Robust Unlabeled Learning

### v1

@@ Abstract
Recent advances in large language models (LLMs) have yielded impressive performance on various tasks, yet they often depend on high-quality feedback that can be costly. Self-refinement methods attempt to leverage LLMs’ internal evaluation mechanisms with minimal human supervision; however, these approaches frequently suffer from inherent biases and overconfidence—especially in domains where the models lack sufficient internal knowledge—resulting in performance degradation. As an initial step toward enhancing self-refinement for broader applications, we introduce an iterative refinement pipeline that employs the Unlabeled-Unlabeled learning framework to improve LLM-generated pseudo-labels for classification tasks. By exploiting two unlabeled datasets with differing positive class ratios, our approach iteratively denoises and refines the initial pseudo-labels, thereby mitigating the adverse effects of internal biases with minimal human supervision. Evaluations on diverse datasets—including low-resource language corpora, patent classifications, and protein structure categorizations—demonstrate that our method consistently outperforms both initial LLM’s classification performance and the self-refinement approaches by cutting-edge models (e.g., GPT-4o and DeepSeek-R1).
Self Iterative Label Refinement via Robust Unlabeled Learning
Hikaru Asano1,2     Tadashi Kozuno3     Yukino Baba1 1The University of Tokyo  2RIKEN AIP  3OMRON SINIC X asano-hikaru19, yukino-baba@g.ecc.u-tokyo.ac.jp, tadashi.kozuno@sinicx.com
@@ 1 Introduction
Rapid advancements in large language models (LLMs) have yielded significant improvements across various downstream tasks and have raised a fundamental research question: How can we further improve an LLM’s capabilities with minimal human supervision? Traditional approaches, such as Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) and its variants (Rafailov et al., 2023; Tang et al., 2024), improve performance through classification tasks (Tang et al., 2024), yet they rely on extensive, costly, and time-consuming human-labeled datasets. Although recent methods like Reinforcement Learning from AI Feedback (RLAIF) (Bai et al., 2022; Lee et al., 2024; Zhu et al., 2024) aim to reduce these costs by replacing human annotations with model-generated signals, their success critically hinges on the reliability of the model’s self-evaluation (Li et al., 2024a; Wang et al., 2024).
Figure 1: Classification accuracy curves over five iterations on three challenging datasets. Ours (Oracle) uses the exact class prior for UU learning, while Ours (few-labeled) estimates these priors from only 50 labeled examples. Our method consistently improves accuracy and outperforms both LLM self-refinement by GPT-4o series and advanced reasoning model DeepSeek-R1
When using an LLM as its own evaluator, it is observed that the model’s inherent biases can harm the reliability of its assessments (Koo et al., 2024; Li et al., 2024b; Liu et al., 2024; Wang et al., 2024), thereby undermining the effectiveness of downstream training (Baumgärtner et al., 2024). While iterative self-refinement (Kim et al., 2023; Madaan et al., 2023) and multi-agent frameworks (Estornell and Liu, 2024) can mitigate simpler biases (e.g., ordering or length biases), LLMs still encounter significant challenges in areas where their internal knowledge is limited (Kamoi et al., 2024a). In such cases, external tools—such as accurate, domain-specific databases—can help address certain blind spots (Wu et al., 2024), but they do not fully eliminate the need for human supervision. Without either robust internal knowledge or dependable huge human input, conventional methods not only struggle to improve performance (Huang et al., 2024; Kamoi et al., 2024a) but also may even experience degradation due to inaccuracies and overconfidence (Huang et al., 2024).
To address these challenges specifically for classification tasks—a crucial first step towards broader self-refinement in LLMs—we introduce an iterative pipeline that refines LLM-generated pseudo-labels using a weakly supervised learning technique known as the Unlabeled-Unlabeled (UU) learning framework (Lu et al., 2019, 2020). Notably, while creating a large, well-maintained labeled dataset requires intensive human supervision, it is relatively straightforward to amass a vast corpus of unlabeled data in the modern era (Sagiroglu and Sinanc, 2013). Motivated by this observation, our approach leverages two unlabeled datasets with differing positive class ratios. Under the simple assumption that one dataset contains a higher proportion of positive examples than the other, our system can effectively learn to distinguish between positive and negative instances without requiring explicit annotations for each example.
@@
We evaluate our approach on several public datasets, including low-resource language corpora, patent classification tasks, and protein structure classification. Notably, as illustrated in Figure 1, even in cases where self-refinement methods based on LLMs—or advanced reasoning models such as DeepSeek-R1 (DeepSeek-AI, 2025)—fail to produce any performance improvement, our iterative UU learning framework successfully refines its outputs and achieves classification performance that surpasses that of both the original LLM and existing self-improvement pipelines. These results underscore the potential of our method not only for specialized text classification but also as a novel approach to achieving self-refinement in LLMs.
In summary, our contributions are threefold:
  • We introduce an iterative pipeline that refines LLM-generated pseudo-labels via UU Learning, reducing noise and boosting classification accuracy with minimal human supervision.
  • We show that the proposed method outperforms both direct LLM classification and self-improvement methods across diverse tasks, including low-resource language, patent, and protein classification.
  • Our approach mitigates annotation noise and offers a scalable, cost-effective solution for high-quality classification in specialized domains with limited expert-labeled data.
@@ 2 Related Work
RLAIF is a popular LLM post-training method (Bai et al., 2022). Its idea to generate feedback by a model itself is called Pseudo-Labeling (PL) and well studied in semi-supervised learning (Scudder, 1965; Lee, 2013). Below, we will review PL and LLM self-training methods related to ours. For a thorough review of each topic, please refer to Yang et al. (2023) and Xiao and Zhu (2025).
@@
PL trains a student model so that its output is close to the output of a teacher model on unlabeled data (Scudder, 1965; Lee, 2013). Various methods for constructing a teacher model has been proposed. For example, $\Pi$-model (Bachman et al., 2014; Sajjadi et al., 2016; Laine and Aila, 2017) and virtual adversarial training (Miyato et al., 2019) perturbs input and/or neural networks of student models. Some methods use weighted average of previous student models’ predictions or weights as a teacher model (Laine and Aila, 2017; Tarvainen and Valpola, 2017), and other methods even try to optimize a teacher model by viewing PL as an optimization problem (Yi and Wu, 2019; Wang and Wu, 2020; Pham et al., 2021). Our work is complimentary to this line of works since what we modify is the risk estimator rather than teacher models.
PL is known to suffer from erroneous labels generated by a model-in-training (Haase-Schutz et al., 2021; Rizve et al., 2021), and this observation well aligns with recent reports on RLAIF that erroneous self-feedback is a major source of failure (Madaan et al., 2023). A straightforward approach is to filter potentially incorrect labels based on confidence (Sohn et al., 2020; Haase-Schutz et al., 2021; Rizve et al., 2021) or the amplitude of loss as in self-paced learning (Bengio et al., 2009; Kumar et al., 2010; Jiang et al., 2015). Another method is refining pseudo-labels in a way similar to label propagation (Zhu and Ghahramani, 2002) but with similarity scores computed using neural networks (Kutt et al., 2024). Our work tackles the issue of erroneous labels by using a risk estimate robust to erroneous self-feedback based on UU learning (Lu et al., 2019, 2020). Even though our approach requires only a minimal change, we observed a significant performance boost.
@@
To enhance the reasoning capabilities of LLMs, early efforts primarily explored various prompt engineering techniques (Brown et al., 2020; Wei et al., 2022; Wang et al., 2023b; Yao et al., 2023). Even with refined prompting strategies, an LLM’s initial outputs can remain limited in some scenarios. Recent studies have proposed iteratively improved answer strategies called self-refinement approaches (Madaan et al., 2023; Kim et al., 2023; Chen et al., 2024b), and our work falls within this lineage.
Self-refinement involves generating responses through agents in distinct roles that iteratively provide feedback (Shinn et al., 2023; Zhu et al., 2023). For example, multi-agent debate frameworks (Estornell and Liu, 2024) use an answering agent and an evaluation agent to collaboratively improve answer quality (Chen et al., 2024a; Du et al., 2024; Smit et al., 2024). However, these methods usually assume that the LLM has enough internal knowledge to generate effective feedback and revisions; when it doesn’t, performance gains can be minimal or even negative (Huang et al., 2024; Li et al., 2024c; Kamoi et al., 2024a, b)—sometimes degrading performance (Huang et al., 2024). Our approach minimizes this reliance: internal knowledge is used only initially for classification, with subsequent improvements relying on extracting features directly from the data via UU learning.
Other work addresses knowledge gaps by retrieving external information or using external tools (Huang et al., 2022; Wang et al., 2023a; Shi et al., 2024; Wu et al., 2024), but setting up such systems can be costly. In contrast, our method requires only a small amount of labeled data for initialization without assuming the availability of external knowledge or tools.
@@ 3.1 Supervised Binary Classification
A classifier $g:\mathbb{R}^{d}\to\mathbb{R}$ outputs a real-valued score, whose sign determines the predicted label. For instance, a neural network can serve as $g$. A loss function :×{±$\ell:\mathbb{R}\times\{\pm 1\}\to[0,\infty)$ then measures how much the prediction disagrees with the true label. Let $R^{+}_{p}(g)=\mathbb{E}_{x\sim p_{p}}[\ell(g(x),+1)]$ denote the loss for true positive data, and $R^{-}_{n}(g)=\mathbb{E}_{x\sim p_{n}}[\ell(g(x),-1)]$ denote the loss for the true negative data. Then, the true risk is expressed as
@@ 3.2 Unlabeled-Unlabeled (UU) Learning
UU learning (Lu et al., 2019) is a technique that allows training a classifier without fully labeled positive and negative datasets, leveraging two unlabeled datasets with different class priors.
@@
$\widetilde{p}_{p}(x)$ $=\theta_{p}\,p_{p}(x)\;+\;\bigl{(}1-\theta_{p}\bigr{)}\,p_{n}(x)$
$\widetilde{p}_{{}_{n}}(x)$ $=\theta_{n}\,p_{p}(x)\;+\;\bigl{(}1-\theta_{n}\bigr{)}\,p_{n}(x).$
@@
$R_{\mathrm{uu}}(g)$ (2)
$\hskip 10.00002pt=aR_{\tilde{p}}^{+}(g)-bR_{\tilde{p}}^{-}(g)-cR_{\tilde{n}}^{+}(g)+dR_{\tilde{n}}^{-}(g),$
@@ 3.3 Robust UU Learning
While UU learning (2) does allow model training without explicit positive/negative labels, comparing the original binary classification risk (1)—which remains nonnegative—against the UU risk (2) shows the UU risk includes negative terms such as $-bR_{\tilde{p}}^{-}(g)$ and $-cR_{\tilde{n}}^{+}(g)$. It has been observed that these negative risk terms can lead to overfitting (Lu et al., 2020).
To mitigate this, Robust UU Learning introduces a generalized Leaky ReLU function $f$ to moderate the reduction of negative risk. Concretely, it normalizes each term of the loss function as (Lu et al., 2020)
$R_{\mathrm{ruu}}(g)$ $=f\left(aR_{\tilde{p}}^{+}(g)-cR_{\tilde{n}}^{+}(g)\right)$
$\hskip 20.00003pt+f\left(dR_{\tilde{n}}^{-}(g)-bR_{\tilde{p}}^{-}(g)\right)$ (3)
where each bracketed term resembles a “normalized” risk under the hypothetical label of being positive or negative, respectively. The function $f$ is given by
$f(x)=\begin{cases}x&\text{if }x>0\\ \lambda x&\text{if }x<0\end{cases}\quad(\lambda<0).$ (4)
Intuitively, $f$ leaves the risk value unchanged when the risk is positive, but for negative risk, it uses $\lambda<0$ to convert it into a positive quantity, thus mitigating the overfitting by negative risk.
@@ 4 Method
Below, we present our iterative framework for combining large language model (LLM) annotations with robust UU learning. The goal is to iteratively refine pseudo-labels generated by the LLM, thus boosting classification accuracy under minimal human supervision. Figure 2 provides a high-level view of this approach, which proceeds in rounds.
@@ 4.1 Overview of the Iterative Pipeline
Our pipeline is composed of three steps:
  1. 1.
    LLM-Based Annotation (Iteration 0). An LLM provides pseudo-labels for an unlabeled corpus (§4.2).
  2. 2.
    Robust UU Learning. We split the corpus into pseudo-positive and pseudo-negative subsets and train a classifier using robust UU learning (§4.3).
  3. 3.
    Re-Labeling. The trained classifier re-labels the entire dataset, producing updated pseudo-positive and pseudo-negative sets for the next iteration (§4.4).
@@ 4.2 Initial Noisy Annotation via LLM
Let $\mathcal{C}=\{x_{1},x_{2},\dots,x_{N}\}$ be a corpus of unlabeled samples. We use the LLM as the initial classifier to assign an initial pseudo-label $\tilde{y}_{i}\in\{+1,-1\}$ to each sample $x_{i}$. Our prompt first provides a concise description of the task, the dataset domain, and the expected answer format (e.g., “Output TRUE or FALSE”). We then give a few-shot examples illustrating how to label an example text, along with a short rationale. Finally, the prompt includes the samples to annotate (see Fig. 6 for the exact prompt).
@@ 4.3 Refinement with Robust UU Learning
Let $\widetilde{\mathcal{C}}_{p}^{(t-1)}$ and $\widetilde{\mathcal{C}}_{n}^{(t-1)}$ denote the pseudo-positive and pseudo-negative sets from iteration $t-1$. Our goal is to train a classifier $g^{(t)}$ (e.g., a neural network) despite noisy labels. To this end, we optimize the robust UU learning objective:
@@ 4.4 Iterative Re-Labeling and Convergence
After training $g^{(t)}$, we re-label the entire dataset:
$\tilde{y}_{i}^{(t)}\;=\;\begin{cases}+1,&\text{if }g^{(t)}(x_{i})>0,\\ -1,&\text{otherwise}.\end{cases}$
We then form updated pseudo-positive and pseudo-negative sets,
$\widetilde{\mathcal{C}}_{p}^{(t)}=\{\,x_{i}\mid\tilde{y}_{i}^{(t)}=+1\}\text{, }\widetilde{\mathcal{C}}_{n}^{(t)}=\{\,x_{i}\mid\tilde{y}_{i}^{(t)}=-1\},$
and use them as input for the next iteration of robust UU learning.
Over several iterations, this process progressively improves the reliability of the pseudo-labels. In the ideal scenario, the positive prior in $\widetilde{\mathcal{C}}_{p}^{(t)}$ converges to 1, and the positive prior in $\widetilde{\mathcal{C}}_{n}^{(t)}$ converges to 0, bringing each corpus ever closer to the gold-standard case of perfectly labeled positive and negative data. When these priors reach 1 and 0, respectively, robust UU learning effectively reduces to standard supervised learning—achieving high accuracy even from initially noisy labels.

### v2

@@ Abstract
Recent advances in large language models (LLMs) have yielded impressive performance on various tasks, yet they often depend on high-quality feedback that can be costly. Self-refinement methods attempt to leverage LLMs’ internal evaluation mechanisms with minimal human supervision; however, these approaches frequently suffer from inherent biases and overconfidence, especially in domains where the models lack sufficient internal knowledge, resulting in performance degradation. As an initial step toward enhancing self-refinement for broader applications, we introduce an iterative refinement pipeline that employs the Unlabeled-Unlabeled learning framework to improve LLM-generated pseudo-labels for classification tasks. By exploiting two unlabeled datasets with differing positive class ratios, our approach iteratively denoises and refines the initial pseudo-labels, thereby mitigating the adverse effects of internal biases with minimal human supervision. Evaluations on diverse datasets, including low-resource language corpora, patent classifications, and protein structure categorizations, demonstrate that our method consistently outperforms both initial LLM’s classification performance and the self-refinement approaches by cutting-edge models (e.g., GPT-4o and DeepSeek-R1). Moreover, we experimentally confirm that our refined classifier facilitates effective post-training alignment for safety in LLMs and demonstrate successful self-refinement in generative tasks as well. [^1] Our code is available at https://github.com/HikaruAsano/self-iterative-label-refinement.
@@ 1 Introduction
Rapid advancements in large language models (LLMs) have yielded significant improvements across various downstream tasks and have raised a fundamental research question: How can we further improve an LLM’s capabilities with minimal human supervision? Traditional approaches, such as Reinforcement Learning from Human Feedback (RLHF) [40] and its variants [42, 53], improve performance through classification tasks [53], yet they rely on extensive, costly, and time-consuming human-labeled datasets. Although recent methods like Reinforcement Learning from AI Feedback (RLAIF) [3, 29, 67] aim to reduce these costs by replacing human annotations with model-generated signals, their success critically hinges on the reliability of the model’s self-evaluation [30, 58].
When using an LLM as its own evaluator, it is observed that the model’s inherent biases can harm the reliability of its assessments [24, 31, 34, 58], thereby undermining the effectiveness of downstream training [4]. While iterative self-refinement [23, 37] and multi-agent frameworks [16] can mitigate simpler biases (e.g., ordering or length biases), LLMs still encounter significant challenges in areas where their internal knowledge is limited [21]. In such cases, external tools, such as accurate, domain-specific databases, can help address certain blind spots [61], but they do not fully eliminate the need for human supervision. Without either robust internal knowledge or dependable huge human input, conventional methods not only struggle to improve performance [18, 21] but also may even experience degradation due to inaccuracies and overconfidence [18].
To address these challenges specifically for classification tasks, a crucial first step towards broader self-refinement in LLMs, we introduce an iterative pipeline that refines LLM-generated pseudo-labels using a weakly supervised learning technique known as the Unlabeled-Unlabeled (UU) learning framework [35, 36]. Notably, while creating a large, well-maintained labeled dataset requires intensive human supervision, it is relatively straightforward to amass a vast corpus of unlabeled data in the modern era [45]. Motivated by this observation, our approach leverages two unlabeled datasets with differing positive class ratios. Under the simple assumption that one dataset contains a higher proportion of positive examples than the other, our system can effectively learn to distinguish between positive and negative instances without requiring explicit annotations for each example.
@@
We evaluate our approach on several public datasets, including low-resource language corpora, patent classification tasks, and protein structure classification. Notably, as illustrated in Figure 3, even in cases where self-refinement methods based on LLMs, or advanced reasoning models such as DeepSeek-R1 [13], fail to produce any performance improvement, our iterative UU learning framework successfully refines its outputs and achieves classification performance that surpasses that of both the original LLM and existing self-improvement pipelines. Moreover, we experimentally show that our refined classifiers, integrated into RLAIF frameworks, effectively achieve safety alignment without extensive human annotation. This underscores our method’s potential for comprehensive, robust LLM self-refinement.
In summary, our contributions are threefold: (i) We introduce an iterative pipeline that refines LLM-generated pseudo-labels via UU Learning, reducing noise and boosting classification accuracy with minimal human supervision; (ii) we demonstrate our method consistently surpasses direct LLM classification and existing self-improvement methods across diverse tasks (e.g., low-resource languages, patents, proteins), enabling scalable, high-quality classification with limited labeled data; and (iii) we experimentally show that our refined classifier facilitates effective post-training alignment for LLMs and highlight its potential for self-refinement in broader generative tasks.
@@ 2 Related Work
RLAIF is a popular LLM post-training method [3]. Its idea to generate feedback by a model itself is called Pseudo-Labeling (PL) and well studied in semi-supervised learning [28, 48]. Below, we will review PL and LLM self-training methods related to ours. For a thorough review of each topic, please refer to Yang et al. [63] and Xiao and Zhu [62].
@@
PL trains a student model so that its output is close to the output of a teacher model on unlabeled data [28, 48]. Various methods for constructing a teacher model has been proposed. For example, $\Pi$-model [2, 27, 46] and virtual adversarial training [39] perturbs input and/or neural networks of student models. Some methods use weighted average of previous student models’ predictions or weights as a teacher model [27, 55], and other methods even try to optimize a teacher model by viewing PL as an optimization problem [41, 57, 65]. Our work is complimentary to this line of works since what we modify is the risk estimator rather than teacher models.
PL is known to suffer from erroneous labels generated by a model-in-training [17, 43], and this observation well aligns with recent reports on RLAIF that erroneous self-feedback is a major source of failure [37]. A straightforward approach is to filter potentially incorrect labels based on confidence [9, 17, 20, 43, 52] or the amplitude of loss as in self-paced learning [6, 20, 25]. Another method is refining pseudo-labels in a way similar to label propagation [68] but with similarity scores computed using neural networks [26]. Our work tackles the issue of erroneous labels by using a risk estimate robust to erroneous self-feedback based on UU learning [35, 36]. Even though our approach requires only a minimal change, we observed a significant performance boost.
@@
To enhance the reasoning capabilities of LLMs, early efforts primarily explored various prompt engineering techniques [8, 59, 60, 64]. Even with refined prompting strategies, an LLM’s initial outputs can remain limited in some scenarios. Recent studies have proposed iteratively improved answer strategies called self-refinement approaches [11, 23, 37], and our work falls within this lineage.
Self-refinement involves generating responses through agents in distinct roles that iteratively provide feedback [50, 69]. For example, multi-agent debate frameworks [16] use an answering agent and an evaluation agent to collaboratively improve answer quality [10, 15, 51]. However, these methods usually assume that the LLM has enough internal knowledge to generate effective feedback and revisions; when it doesn’t, performance gains can be minimal or even negative [18, 21, 22, 32], sometimes degrading performance [18]. Our approach minimizes this reliance: internal knowledge is used only initially for classification, with subsequent improvements relying on extracting features directly from the data via UU learning.
Other work addresses knowledge gaps by retrieving external data or using external tools [19, 49, 56, 61], though such setups can be costly. In contrast, our method needs only a small amount of labeled data for initialization, without relying on external resources.
@@ 3.1 Supervised Binary Classification
A classifier $g:\mathbb{R}^{d}\to\mathbb{R}$ outputs a real-valued score, whose sign determines the predicted label. For instance, a neural network can serve as $g$. A loss function $\ell:\mathbb{R}\times\{\pm 1\}\to[0,\infty)$ then measures how much the prediction disagrees with the true label. Let $R^{+}_{p}(g)=\mathbb{E}_{x\sim p_{p}}[\ell(g(x),+1)]$ denote the loss for true positive data, and $R^{-}_{n}(g)=\mathbb{E}_{x\sim p_{n}}[\ell(g(x),-1)]$ denote the loss for the true negative data. Then, the true risk is expressed as
@@ 3.2 Unlabeled-Unlabeled (UU) Learning
UU learning [35] is a technique that allows training a classifier without fully labeled positive and negative datasets, leveraging two unlabeled datasets with different class priors.
@@
$\widetilde{p}_{p}(x)=\theta_{p}\,p_{p}(x)+\bigl(1-\theta_{p}\bigr)\,p_{n}(x),\quad\widetilde{p}_{n}(x)=\theta_{n}\,p_{p}(x)+\bigl(1-\theta_{n}\bigr)\,p_{n}(x).$
@@
$R_{\mathrm{uu}}(g)=aR_{\tilde{p}}^{+}(g)-bR_{\tilde{p}}^{-}(g)-cR_{\tilde{n}}^{+}(g)+dR_{\tilde{n}}^{-}(g),$ (2)
@@ 3.3 Robust UU Learning
While UU learning (2) does allow model training without explicit positive/negative labels, comparing the original binary classification risk (1), which remains nonnegative, against the UU risk (2) shows the UU risk includes negative terms such as $bR_{\tilde{p}}^{-}(g)$ and $cR_{\tilde{n}}^{+}(g)$. It has been observed that these negative risk terms can lead to overfitting [36].
To mitigate this, Robust UU Learning introduces a generalized Leaky ReLU function $f$ to moderate the reduction of negative risk. Concretely, it normalizes each term of the loss function as [36]
$R_{\mathrm{ruu}}(g)=f\left(aR_{\tilde{p}}^{+}(g)-cR_{\tilde{n}}^{+}(g)\right)+f\left(dR_{\tilde{n}}^{-}(g)-bR_{\tilde{p}}^{-}(g)\right)$ (3)
where each bracketed term resembles a “normalized” risk under the hypothetical label of being positive or negative, respectively. The function $f$ is defined as $f(x)=x$ if $x>0$, and $f(x)=\lambda x$ if $x<0$, where $\lambda<0$. Intuitively, $f$ preserves positive risk values while converting negative risk into positive values using $\lambda<0$, thereby reducing overfitting caused by negative risk terms.
@@ 4 Method
Below, we present our iterative framework for combining LLM annotations with robust UU learning. The goal is to iteratively refine pseudo-labels generated by the LLM, thus boosting classification accuracy under minimal human supervision. Figure 1 provides an overview of this pipeline.
@@ 4.1 Overview of the Iterative Pipeline
Our pipeline proceeds in three steps, (i) LLM-Based Annotation (Iteration 0): an LLM provides pseudo-labels for an unlabeled corpus (§4.2); (ii) Robust UU Learning: we split the corpus into pseudo-positive and pseudo-negative subsets and train a classifier via robust UU learning (§4.3); and (iii) Re-Labeling: the trained classifier re-labels the entire dataset, producing refreshed pseudo-positive and pseudo-negative sets for the next iteration (§4.4).
@@ 4.2 Initial Noisy Annotation via LLM
Let $\mathcal{C}=\{x_{1},x_{2},\dots,x_{N}\}$ be a corpus of unlabeled samples. We use the LLM as the initial classifier to assign an initial pseudo-label $\tilde{y}_{i}\in\{+1,-1\}$ to each sample $x_{i}$. Our prompt first provides a concise description of the task, the dataset domain, and the expected answer format (e.g., “Output TRUE or FALSE”). We then give a few-shot examples illustrating how to label an example text, along with a short rationale. Finally, the prompt includes the samples to annotate (see Figure 5 for the exact prompt).
@@ 4.3 Refinement with Robust UU Learning
Let $\widetilde{\mathcal{C}}_{p}^{(t-1)}$ and $\widetilde{\mathcal{C}}_{n}^{(t-1)}$ denote the pseudo-positive and pseudo-negative sets from iteration t1t-1. Our goal is to train a classifier $g^{(t)}$ (e.g., a neural network) despite noisy labels. To this end, we optimize the robust UU learning objective:
@@ 4.4 Iterative Re-Labeling and Convergence
After training $g^{(t)}$, we re-label the entire dataset: $\tilde{y}_{i}^{(t)}:=\operatorname{sign}\bigl(g^{(t)}(x_{i})\bigr)\in\{+1,-1\}$. From these labels, we form updated pseudo-positive $\widetilde{\mathcal{C}}_{p}^{(t)}=\{\,x_{i}\;\mid\;\tilde{y}_{i}^{(t)}=+1\,\}$ and pseudo-negative $\widetilde{\mathcal{C}}_{n}^{(t)}=\{\,x_{i}\;\mid\;\tilde{y}_{i}^{(t)}=-1\,\}$ sets, used in the next robust UU learning iteration.
Over several iterations, this process progressively improves the reliability of the pseudo-labels. In the ideal scenario, the positive prior in $\widetilde{\mathcal{C}}_{p}^{(t)}$ converges to 1, and the positive prior in $\widetilde{\mathcal{C}}_{n}^{(t)}$ converges to 0, bringing each corpus ever closer to the gold-standard case of perfectly labeled positive and negative data. When these priors reach 1 and 0, respectively, robust UU learning effectively reduces to standard supervised learning, achieving high accuracy even from initially noisy labels.